Circular No. 6221

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995Y IN NGC 410 J. Mueller reports her discovery of an apparent supernova (mag ∼ 18)

located 7′′ east and 25′′ south of the center of NGC 410 (α = 1^{h}11^{m}.0, δ = +33^{o}09′, equinox 2000.0); the IV-N discovery plate was taken on Aug. 28 UT by J. D. Mendenhall and Mueller with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. M. Schmidt, D. Schneider, and R. Lucinio obtained a spectrogram of SN 1995Y with the 5-m Hale Telescope (+ 4-shooter + spectrograph) on Aug. 31, and this indicates that the supernova is of type-I and is probably several months past maximum.

SUPERNOVA 1995Z IN UGC 937 Mueller also reports her discovery of an apparent supernova (mag ∼

18.5) located 11′′ east and 5′′ north of the center of UGC 937 (α = 1^{h}23^{m}.6, δ = +32^{o}39′, equinox 2000.0); the IV-N discovery plate was taken on Aug. 28 UT by J. D. Mendenhall and Mueller with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. N. Reid notes that no object was present at this position on a IV-N Survey plate taken on 1989 Dec. 4 by C. Brewer and Mendenhall. Schmidt and Schneider obtained a spectrogram (as above) of SN 1995Z on Sept. 1, and this shows strong Hα emission that, when coupled with Reid’s observation, suggests a type-II supernova.

NOVA CASSIOPEIAE 1995 C. E. Woodward, D. Harker, and J. Smilowitz, University of Wyoming;

and R. Rudy and R. Kellogg, The Aerospace Corporation, report the following near-infrared magnitudes (with 1σ errors of 0.09 mag) for N Cas 1995, obtained on Aug. 29.3 UT using the 2.31-m telescope of the Wyoming Infrared Observatory (+ Aerospace near-infrared camera + NICMOS3 array): J = 7.89; H = 7.65; [2.00–2.32 μm] = 7.49.

COMET 41P/TUTTLE-GIACOBINI-KRESÁK Total visual magnitude estimates (cf. IAUC 6210): Aug. 25.16 UT, 9.0

(C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 26.86, 9.3 (J. Carvajal, Avila, Spain, 0.44-m reflector); 31.16, 10.0: (Morris).

1995 September 1 Daniel W. E. Green